# 0_home_mhaseliu_Documents_Notes_Research_Multiple_Fourier_Mult_Func_Eq.bib

### v1

@@ 1   Introduction
Let $M$ be a symmetric $m\times m$ matrix with coefficients in $\mathbb{Z}/n\mathbb{Z}$. Define $M^{\prime}$ to be another symmetric $m\times m$ matrix such that
@@
$=\sum_{f_{1},\ldots,f_{s}\in\mathbb{F}_{q}[t]^{+}}\sum_{d_{s+1},\ldots,d_{m}}b\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{f_{s+1}\in\mathcal{M}_{d_{s+1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M^{\prime}\right)u_{1}^{d_{1}}\cdots u_{m}^{d_{m}},$
@@
The assumption on $n$ is mainly due to the fact that the unique multiplicative character $\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ of order two is simply $\chi^{n/2}$. Note that if $\chi$ has odd order and $q$ is not a power of 2, we can replace $\chi$ with its square root, double $n$, and double the entries of $M$ to get an equivalent series where $n$ is even.
@@ 2   Preliminaries
The third subsection is about proving some (probably weak) bounds on counting different kinds of trees.
The fourth subsection gives bounds for the cohomology of arbitrary lisse sheaves on the Kontsevich moduli space ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ (quotiented by a Young subgroup). The main input is the bounds on tree counts from the previous subsection.
@@ 2.1 Notation and Sawin’s construction
    If $X$ is a variety, $j\colon U\subset X$ is the inclusion of an open subset, and $A$ is a perverse sheaf on $U$, then there is a perverse sheaf $j_{!*}A$ on $X$, defined as the unique extension of $A$ that has no non-trivial sub-objects or quotients supported on $X\backslash U$.
@@ 2.3 Elementary combinatorial lemmas on trees
For such a tree $T$ and a non-leaf vertex $v$, consider the $d(v)$ (i.e. the valency) neighbors of $v$, including the leaves. If we remove the edge connecting $v$ to a neighbor $w$, then the connected component containing $w$ can be viewed as a rooted tree $T_{w}$ with some number of labels. Each such neighbor defines a new rooted sub-tree with the neighbor as a root. Let $\lambda(v)$ be a partition of $d(v)$ and $S_{\lambda(v)}$ be the product of symmetric groups $S_{\lambda(v)_{1}}\times S_{\lambda(v)_{2}}\times\cdots$ defined as follows:
@@
    There are at most $r+1$ non-leaf vertices.
@@
    The sum of the valences over all non-leaf vertices is at most $3r$.
@@
    The sum of the valences over all non-leaf vertices minus the number of non-leaf, non-root vertices is at most $2r$.
@@
By assumption, the left-hand side is at least $3s$, which means $s\leq r$, i.e. there are at most $r+1$ non-leaf vertices. This also implies $d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)\leq 3r$ and $d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)-s=r+s\leq 2r$.
Finally, suppose for the sake of contradiction that a non-leaf vertex $v$ has valency at least $r+2$. Also, suppose $v$ has $\ell$ leaves, which is at most $r$. Then, at least $r+2-\ell-1\geq 1$ of these neighbors $w$ are not leaves or the root. By the valency assumption, each $T_{w}$ necessarily contains at least two leaves. Then, $2\left(r+2-\ell-1\right)+\ell$ is at most the total number of leaves, which is $r$. But this implies $\ell\geq r+2$, which is clearly absurd. ∎
@@
We can think of any tree in $\mathcal{T}^{\prime}$ as a rooted unlabelled tree with at least $r+1$ vertices and at most $r+1+r=2r+1$ vertices by lemma 13.
@@
(${r\choose d_{1},\ldots,d_{m}}\left(C_{r}+\cdots+C_{2r}\right)\leq\frac{r!}{\left(\left(r/m\right)!\right)^{m}}4^{2r}.$
@@
(${r\choose d_{1},\ldots,d_{m}}\left(C_{r}+\cdots+C_{2r}\right)$ $\leq\frac{\sqrt{2\pi r}\left(\frac{r}{e}\right)^{r}e^{\frac{1}{12r}}}{\left(\sqrt{2\pi\frac{r}{m}}\left(\frac{r/m}{e}\right)^{r/m}e^{\frac{1}{12\frac{r}{m}+1}}\right)^{m}}16^{r}$
$=\frac{m^{m/2}e^{\frac{1}{12r}-\frac{m^{2}}{12r+m}}}{\left(\sqrt{2\pi r}\right)^{m-1}}\left(16m\right)^{r}$
@@ 2.4 Bounds for the cohomology of lisse sheaves on the moduli space of stable maps
By lemma 12, the sum of valences of non-leaf vertices minus the number of non-leaf, non-root vertices is at most $2r$. As such, we have
@@
$\leq 2^{2r}\sum_{T\in\mathcal{T}}\left({d(\text{root})\choose\lambda(\text{root})}\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash 1}\right).$
@@ 3   Relationship between $a$-coefficients and Fourier transforms
Let us continue to use the notation from subsection 2.1. Moreover, as in the hypotheses of theorem 6, we assume
@@ 6   Meromorphic continuation of multiple Dirichlet series
Indeed, observe that showing convergence for

### v2

# Mult_Func_Eq.bib

@@ 1   Introduction
Let $M$ be a symmetric $m\times m$ matrix with coefficients in $\mathbb{Z}/n\mathbb{Z}$ for $n$ even. Define $M^{\prime}$ to be another symmetric $m\times m$ matrix such that
@@
$\coloneqq\sum_{f_{1},\ldots,f_{s}\in\mathbb{F}_{q}[t]^{+}}\sum_{d_{s+1},\ldots,d_{m}}b\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{f_{s+1}\in\mathcal{M}_{d_{s+1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M^{\prime}\right)u_{1}^{d_{1}}\cdots u_{m}^{d_{m}},$
@@
The assumption on the parity of $n$ is mainly due to the fact that the unique multiplicative character $\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ of order two is simply $\chi^{n/2}$. Note that if $\chi$ has odd order and $q$ is not a power of 2, we can replace $\chi$ with its square root, double $n$, and double the entries of $M$ to get an equivalent series where $n$ is even.
@@ 2   Preliminaries
The third subsection is about proving bounds on the counts of different kinds of trees.
The fourth subsection gives bounds for the cohomology of arbitrary lisse sheaves on the Kontsevich moduli space ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ quotiented by a Young subgroup. The main input is the bounds on tree counts from the previous subsection.
@@ 2.1 Notation and Sawin’s construction
    If $X$ is a variety, $j\colon U\subset X$ is the inclusion of an open subset, and $A$ is a perverse sheaf on $U$, then there is a perverse sheaf $j_{!*}A$ on $X$, called the intermediate extension of $A$, defined as the unique extension of $A$ that has no non-trivial sub-objects or quotients supported on $X\backslash U$.
@@ 2.3 Elementary combinatorial lemmas on trees
For such a tree $T$ and a non-leaf vertex $v$, consider the $d(v)$ (i.e. the valency) neighbors of $v$, including the leaves. If we remove the edge connecting $v$ to a neighbor $w$, then the connected component containing $w$ can be viewed as a rooted tree $T_{w}$ with some number of labels. Each such neighbor defines a new rooted sub-tree. Let $\lambda(v)$ be a partition of $d(v)$ and $S_{\lambda(v)}$ be the product of symmetric groups $S_{\lambda(v)_{1}}\times S_{\lambda(v)_{2}}\times\cdots$ defined as follows:
@@
    There are at most $r$ non-leaf vertices.
@@
    The sum of the valences over all non-leaf vertices is at most $23r-$ 2.
@@
    The sum of the valences over all non-leaf vertices minus the number of non-leaf, non-root vertices is at most $12r-$ 1.
@@
By assumption, the left-hand side is at least $13s+$ 1, which means $s\leq r-1$, i.e. there are at most $r$ non-leaf vertices. This also implies $d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)\leq 3r-2$ and $d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)-s=r+s\leq 2r-1$.
Finally, suppose for the sake of contradiction that a non-leaf vertex $v$ has valency at least $r+2$. Also, suppose $v$ has $\ell$ leaves, which is at most $r$. Then, at least $r+2-\ell-1\geq 1$ of these neighbors $w$ are not leaves or the parent of $v$. By the valency assumption, each $T_{w}$ necessarily contains at least two leaves. Then, $2\left(r+2-\ell-1\right)+\ell$ is at most the total number of leaves, which is $r$. But this implies $\ell\geq r+2$, which is clearly absurd. ∎
@@
We can think of any tree in $\mathcal{T}^{\prime}$ as a rooted unlabelled tree with at least $r+1$ vertices and at most $r+r=2r$ vertices by lemma 13.
@@
(${r\choose d_{1},\ldots,d_{m}}\left(C_{r}+\cdots+C_{2r-1}\right)\leq\frac{r!}{\left(\left(r/m\right)!\right)^{m}}4^{2r-1}.$
@@
(${r\choose d_{1},\ldots,d_{m}}\left(C_{r}+\cdots+C_{2r}\right)$ $\ll_{m}\frac{\sqrt{2\pi r}\left(\frac{r}{e}\right)^{r}e^{\frac{1}{12r}}}{\left(\sqrt{2\pi\frac{r}{m}}\left(\frac{r/m}{e}\right)^{r/m}e^{\frac{1}{12\frac{r}{m}+1}}\right)^{m}}16^{r}$
$\ll_{m}\frac{m^{m/2}e^{\frac{1}{12r}-\frac{m^{2}}{12r+m}}}{\left(\sqrt{2\pi r}\right)^{m-1}}\left(16m\right)^{r}$
@@ 2.4 Bounds for the cohomology of lisse sheaves on the moduli space of stable maps
By lemma 12, the sum of valences of non-leaf vertices minus the number of non-leaf, non-root vertices is at most $12r-$ 1. As such, we have
@@
$\leq 2^{2r-1}\sum_{T\in\mathcal{T}}\left({d(\text{root})\choose\lambda(\text{root})}\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash 1}\right).$
@@ 3   Relationship between $a$-coefficients and Fourier transforms
Let us continue to use the notation from subsection 2.1. Moreover, as in the hypotheses of theorem 6, we assume for this section that
@@ 6   Meromorphic continuation of multiple Dirichlet series
Indeed, observe that srhowing convergence for